the ratio of the area of the regular polygon of n sides circumscribed about a circle to the area of the regular polygon of equal no. of sides inscribed in the circle is 4:3.find the value of n 

Grade:10

1 Answers

Aman Bansal
12 years ago

Dear Sathyaram,

Let the polygon circumscribed about a circle be denoted by  PQRS........          & 

the polygon inscribed in the circle be    ABCD.........     and their centres be   O .

in polygon   PQRS........   let the mid-point of PQ be Y so OY is the radius of the circle.

 i recommend you to take pen and paper and do as directed .

let side PQ = x

angle POQ = 360/n     therefore   angle POY = 180/n

tan 180/n = x/2r                                     where r is the radius of the circle

x = 2r tan 180/n

area of triangle POQ = xr/2

area of triangle POQ = (2r2)/2 tan 180/n

area of polygon PQRS............= n (area of triangle POQ)                                              (1)-

in polygon ABCD..............

OB is the radius of the circle  , D is the mid-point of AB and angle AOB= 360/n

angle BOD = 180/n

sin 180/n = y/2r                                      where y is the length AB

y = 2r sin 180/n

similarly length OD = r cos 180/n

area of triangle AOB = y(OD)/2

area of triangle AOB = 2r2/2 (sin 180/n)( cos 180/n)

area of polygon ABCD........... = n(area of triangle AOB)

(area of polygon PQRS............)/(area of polygon ABCD...........) = 4/3

substituting the values we will get  a equation as

cos 180/n = √3/2

cos 180/n = cos 30

180/n = 30

 

therefore n =6

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert